# A Comprehensive Graph Pooling Benchmark: Effectiveness, Robustness and Generalizability

### v1

@@ C.1 Graph Classification
The graph classification model utilized in this study shares an identical structure across 15 different pooling methods. Primarily, the model comprises three SAGEConv layers with ReLU activation functions and two pooling layers, followed by a global average pooling layer. Both the hidden and output channels are set to 64. Additionally, the embedding output from the global average pooling layer is processed through a linear layer with ReLU activation and dimensions (64, 32), and subsequently through another linear layer without any activation function and dimensions (32, number of classes). The final output is obtained by applying a softmax function to the embedding output. Due to constraints in computational resources and time, hyperparameter tuning was not performed. All models employed the Adam optimizer with a learning rate set to 0.001 and were trained for 200 epochs using the negative log-likelihood loss function. The data was uniformly split into training, validation, and test sets with ratios of 70%, 15%, and 15%, respectively. Each experiment was run five times with different random seeds.
@@ C.2 Graph Regression
For graph regression, we adopted the backbone network inspired by MESPool [92]. Primarily, the model comprises three GINConv layers with ReLU activation functions and BatchNorm, along with two pooling layers, followed by a global average pooling layer. Both the hidden and output channels are set to 64. Additionally, the embedding output from the global average pooling layer is processed through a linear layer with ReLU activation and dimensions (64, 32). Due to constraints in computational resources and time, hyperparameter tuning was not performed. All models employed the Adam optimizer with a learning rate set to 0.001 and were trained for 200 epochs using the negative log-likelihood loss function. The data was uniformly split into training, validation, and test sets with ratios of 80%, 10%, and 10%, respectively. Each experiment was run five times with different random seeds.
@@ D.1 Graph Classification
Table A.1 provides descriptive statistics of the selected datasets, revealing that our chosen datasets encompass graph data of varying scales and features. This diversity establishes a robust foundation for benchmarking. The following are detailed descriptions of these datasets:
@@
PROTEINS is a collection of graphs representing protein structures where nodes denote secondary structure elements (SSEs) and edges indicate neighborhood relationships between these SSEs. The primary aim of this dataset is to facilitate the classification of proteins into different structural classes based on their amino acid sequences and structural characteristics. Each graph in the dataset is labeled according to its protein class, and the dataset encompasses a diverse range of protein structures, making it a valuable resource for studying the application of graph-based learning methods in bioinformatics and structural biology [52].
PROTEINS_full is an extended version of the PROTEINS dataset, consisting of protein structure graphs for the task of graph classification. Each graph represents a protein, where nodes correspond to secondary structure elements (SSEs) such as alpha helices and beta sheets, and edges represent neighborhood relationships between these SSEs based on spatial proximity or sequential adjacency [52].
@@
NCI1 is a collection of chemical compound graphs derived from the National Cancer Institute (NCI) database, specifically used for the task of graph classification. Each graph in the dataset represents a chemical compound, where nodes correspond to atoms and edges represent the bonds between them. The primary objective of this dataset is to facilitate the classification of compounds based on their ability to inhibit or interact with cancer cell lines, thus aiding in drug discovery and cancer research. The dataset is labeled with two classes, indicating the compounds’ biological activity against a specific cancer cell line [51].
NCI109 is a collection of chemical compound graphs, derived from the National Cancer Institute (NCI) database, aimed at graph classification tasks. Each graph in the dataset represents a chemical compound, with nodes corresponding to atoms and edges representing the bonds between them. The dataset is specifically labeled to reflect the compounds’ bioactivity against cancer cell lines, similar to its counterpart NCI1. NCI109 includes two classes, distinguishing between compounds based on their ability to inhibit or interact with a particular cancer cell line [51].
ENZYMES is a collection of protein graphs specifically designed for the classification of enzymes into one of six EC (Enzyme Commission) top-level classes. Each graph represents a protein, where nodes denote secondary structure elements (SSEs) and edges indicate spatial or sequential neighborhood relationships between these SSEs. The dataset includes attributes for each node, reflecting the physicochemical properties of the amino acids they represent [54].
@@
BZR is a collection of molecular graphs representing chemical compounds, specifically designed for the task of graph classification. Each graph in the dataset corresponds to a chemical compound where nodes represent atoms and edges denote the chemical bonds between them. The BZR dataset is focused on classifying compounds based on their biological activity related to binding with benzodiazepine receptors, which are significant in pharmacology for their role in the effects of various drugs on the central nervous system. Each graph is labeled to indicate whether the compound binds to the benzodiazepine receptor, making this dataset valuable for studying the interaction between chemical compounds and biological targets [50].
@@
IMDB-M is a collection of social network graphs derived from the Internet Movie Database (IMDB). Each graph represents a collaboration network from movies, where nodes correspond to actors or actresses, and edges indicate that the two actors appeared in the same movie. The dataset is specifically designed for the task of multi-class graph classification, with labels indicating the genre of the movie (e.g., Action, Comedy, Drama). This dataset comprises three classes, reflecting the movie genres, and aims to facilitate the classification of social network structures based on their topological features [55].
COLORS-3 is a collection of artificially generated graphs designed for the task of graph classification. Each graph in the dataset represents a structure where nodes are assigned one of three possible colors. The dataset’s primary challenge is to classify these graphs based on their color distribution patterns and structural properties. COLORS-3 provides a controlled environment to test and benchmark graph-based learning algorithms, particularly focusing on the ability to distinguish graphs with different color configurations [56].
@@
D&D is a collection of protein structure graphs designed for the task of graph classification. Each graph in this dataset represents a protein, with nodes corresponding to amino acids and edges representing the spatial or sequential proximity between these amino acids. The primary objective of the D&D dataset is to classify proteins into one of two categories: enzymes or non-enzymes [53].
Table A.1: Summary statistics of datasets for graph classification.
@@ D.2 Graph Regression
Table A.2 provides an overview of the selected datasets in terms of their tasks, compounds and their features, recommended splits, and metrics. A more detailed description is provided below.
QM8 is a benchmark dataset in computational chemistry, designed to facilitate the development and evaluation of machine learning models for quantum mechanical property prediction. It contains approximately 21,786 molecular structures, each characterized by their calculated properties using quantum chemistry methods, specifically focusing on electronic spectra [8, 94].
BACE is a collection of biochemical data used to evaluate computational methods for drug discovery. The dataset includes a total of 1,522 compounds, each annotated with their binding affinities, as well as molecular descriptors and fingerprints to facilitate the development and assessment of machine learning modelsa [8, 95].
ESOL is a prominent resource in cheminformatics, designed for evaluating machine learning models on the prediction of aqueous solubility of small molecules. The dataset, derived from the work of Delaney, encompasses a diverse range of chemical compounds with experimentally determined solubility values expressed in logS, where S is the solubility in mols per liter. It includes 1128 compounds, serving as a benchmark for solubility prediction tasks [96, 8].
FreeSolv is a collection of experimental and calculated hydration free energies for small molecules in aqueous solution. It comprises data for a wide range of organic molecules, providing both experimental values and calculated predictions based on molecular dynamics simulations [97, 8].
@@
Table A.2: Details of datasets for graph regression.
@@ D.3 Node Classification
Table A.3 presents descriptive statistics of the seven datasets used for node classification. It is evident that there is a significant variance in the scale of the selected datasets, each possessing distinct characteristics. Further background information and details about these datasets are provided below.
Cora is a benchmark commonly utilized in the evaluation of graph-based machine learning algorithms, comprises a collection of 2,708 scientific publications classified into seven distinct categories. These categories encompass a wide range of research topics in the field of machine learning. Each publication in the dataset is represented as a node in a citation network, where edges indicate citation relationships between papers [57].
CiteSeer is a widely used citation network dataset. It comprises scientific publications categorized into six classes, with each publication represented by a 3,327-dimensional binary vector indicating the presence or absence of specific words. Additionally, the dataset includes 4,552 citation links among these publications, forming a directed graph where each node represents a paper, and each edge denotes a citation from one paper to another [57].
PubMed consists of scientific publications from the PubMed database, categorized into three classes based on their Medical Subject Headings (MeSH) terms. Each publication is represented as a node, and the citation relationships between the publications form the edges of the graph. Additionally, each node is characterized by a sparse bag-of-words feature vector derived from the content of the corresponding publication [57].
Table A.3: Summary statistics of datasets for node classification.
@@
Cornell, Texas, and Wisconsin originate from the citation networks of their respective universities and consist of nodes representing web pages and edges representing hyperlinks between these pages. Each node is labeled with a class, typically corresponding to the topic of the web page, allowing for tasks such as node classification and link prediction. The datasets vary in size, with Cornell, Texas, and Wisconsin having 183, 183, and 251 nodes respectively, and are characterized by relatively small and sparse networks compared to larger citation datasets [58].
Github includes rich node attributes representing the diverse features of developers, such as their interests, skills, and contributions to various repositories. Additionally, the edges within the network capture the interactions and collaborations among developers, creating a multi-faceted graph structure [59].
@@ D.4 Out-of-distribution shifts
Size shifts For the selected datasets, including D&D, PROTEINS, NCI1, and NCI109, we utilized the data provided by the authors of size-invariant-GNNs [64]. In this setup, the graphs with the smallest 50% of nodes are used as the training set, those with the largest 20% of nodes are used as the test set, and the remaining graphs were used as the validation set. The data can be downloaded from https://www.dropbox.com/s/38eg3twe4dd1hbt/data.zip.
Density shifts For the selected datasets, we divide the datasets based on graph density: the 50% of graphs with the lowest density are used as the training set, the 20% with the highest density are used as the test set, and the remaining graphs are used as the validation set. After applying density shifts, the following densities are observed: for D&D, the training set density is 0.0274, the validation set density is 0.0567, and the test set density is 0.1245; for PROTEINS, the training set density is 0.1709, the validation set density is 0.4183, and the test set density is 1.0752; for NCI1, the training set density is 0.1229, the validation set density is 0.1976, and the test set density is 0.2922; for NCI109, the training set density is 0.1248, the validation set density is 0.2000, and the test set density is 0.2916.
@@ Appendix E Additional Experiments
Table A.4 presents the results of fixing the backbone as GraphConv and setting the pooling ratios of the sparse pooling method to 0.1, 0.3, 0.5, 0.7, and 0.9, respectively. The pooling ratios for dense pooling remain unchanged. As shown in Table A.4, with the increase in pooling ratio, accuracy exhibits a fluctuating upward trend. Among the sparse pooling methods, HGPSLPool and KMISPool consistently demonstrate superior performance. However, changes in pooling ratio also result in varying stability across different methods. Furthermore, compared to binary classification tasks, the accuracy improvement with increasing pooling ratio is more pronounced in multi-class classification tasks. This suggests that researchers should consider using larger pooling ratios when conducting studies on multi-class classification.
@@
Table A.5 presents the results of using GCNConv as the backbone model while varying the pooling ratio. It can be observed that the overall accuracy of GCNConv is lower than that of GraphConv. Generally, as the pooling ratio increases, classification accuracy improves, with dense pooling methods proving to be more powerful than sparse pooling methods, albeit at the cost of increased computation time. In GCNConv, among the sparse pooling methods, CGIPool and KMISPool show the best performance. Additionally, the impact of pooling ratio on multi-class classification accuracy remains significantly greater than its impact on binary classification accuracy.
@@
Table A.6 presents the results of varying the pooling ratio (0.1, 0.3, 0.5, 0.7, 0.9) in node classification tasks using GraphConv as the backbone model. Overall, the change in pooling ratio has a minimal impact on the accuracy of node classification tasks. For most methods, accuracy does not improve with an increased pooling ratio but rather fluctuates. The two best-performing methods overall are TopKPool and GSAPool. Additionally, higher pooling ratios result in increased computational costs for some methods, such as HGPSLPool and ASAPool. This suggests that when performing node classification, graph pooling methods should be used more cautiously, and excessive concern over pooling ratios may be unnecessary.
@@
Table A.7 presents the results of node classification with GCNConv as the backbone model, varying the pooling ratio from 0.1, 0.3, 0.5, 0.7, to 0.9. It can be observed that, except for HGPSLPool, the performance of most methods with GCNConv is inferior to that with GraphConv. The performance of different pooling methods varies across different pooling ratios, making it difficult to identify a single pooling method that consistently outperforms others at all pooling ratios. Nevertheless, CGIPool and KMISPool still relatively demonstrate the best overall performance. Additionally, increasing the pooling ratio does not consistently improve node classification tasks. Lastly, GCNConv has higher memory consumption compared to GraphConv, particularly for ASAPool, PANPool, COPool, and HGPSLPool.
@@
In summary, based on the aforementioned experiments, GraphConv is recommended as the backbone model over GCNConv for both graph classification and node classification tasks due to its higher accuracy and lower computational resource consumption. For graph classification tasks, selecting a larger pooling ratio generally improves results, especially for multi-class classification tasks. In contrast, for node classification tasks, the choice of pooling ratio is not very important.
@@
Table A.4: Results of graph classification using GraphConv as backbone for different graph pooling methods and pooling ratios. $\pm$ denotes standard deviation. OOM denotes out of GPU memory. OOT denotes cannot be computed within 12 hours.
@@
Table A.5: Results of graph classification using GCNConv as backbone for different graph pooling methods and pooling ratios. $\pm$ denotes standard deviation. OOM denotes out of GPU memory. OOT denotes cannot be computed within 12 hours.
@@
Table A.6: Results of node classification using GraphConv as backbone for different graph pooling methods and pooling ratios. $\pm$ denotes standard deviation. OOM denotes out of GPU memory. OOT denotes cannot be computed within 12 hours.
@@
Table A.7: Results of node classification using GCNConv as backbone for different graph pooling methods and pooling ratios. $\pm$ denotes standard deviation. OOM denotes out of GPU memory. OOT denotes cannot be computed within 12 hours.

### v2

@@ C.1 Graph Classification
The classification model comprises three primary components: SAGEConv layers, pooling methods, and a global average pooling layer. The hidden and output channels for this model are both set to 64. Initially, the data passes through three SAGEConv layers with ReLU activation functions, followed by two pooling layers as well as a global average pooling layer. The embedding output from this global layer undergoes further processing through a linear layer with ReLU activation, with dimensions (64, 32), followed by another linear layer without any activation function but with dimensions (32, number of classes). The final output can be available after applying softmax to the embedding output. Hyperparameter tuning is not conducted due to limitations in computational resources and time. All models use the Adam optimizer with a learning rate of 0.001 and are trained for 200 epochs by minimizing the negative log-likelihood loss function. The whole data is divided into training, validation, and test sets (70%, 15%, and 15% ratios), and each trial is repeated five times with different random seeds.
@@ C.2 Graph Regression
We use a backbone network inspired by MESPool [85] for graph regression. The model mainly consists of three GINConv layers with ReLU activation functions and BatchNorm, along with two pooling layers, followed by a global average pooling layer. All channels (both hidden and output) are set to 64. The embedding output from the global average pooling layer passes through another linear layer with ReLU activation, having dimensions (64, 32). Hyperparameter tuning is not conducted due to limitations in computational resources and time. All models use the Adam optimizer with a learning rate of 0.001 and are trained for 200 epochs by minimizing the negative log-likelihood loss function. The data is split uniformly into training, validation, and test sets in ratios of 80%, 10%, and 10% respectively. Each experiment is carried out five times using different random seeds.
@@ D.1 Graph Classification
Table 7 provides descriptive statistics of the selected datasets, which indicates that our chosen datasets encompass graph data of varying scales and features. This diversity establishes a robust foundation for our benchmarking research. The following are detailed descriptions of these datasets:
@@
PROTEINS represents protein structures; nodes denote secondary structure elements (SSEs) and the edges indicate the relationships between these SSEs that are in close proximity. The primary goal is to classify each protein into different classes, i.e., a non-enzyme or an enzyme based on their amino acid sequences and topological characteristics. The dataset covers diverse protein structures [54].
PROTEINS_full is an extended version of PROTEINS, which includes the node attributes as well. The node attribute corresponds to different properties of SSEs including the polarity and the polarizability [54].
@@
NCI1 is a collection of chemical compound graphs. Originating from the National Cancer Institute (NCI) database, each graph sample is a compound in which nodes represent atoms and edges represent the bonds between them. The dataset is binary-class labeled, indicating biological activity via compounds’ anti-cancer activity against specific cell lines [53].
NCI109 is also a collection of chemical compound graphs derived from the NCI database. Similarly, each node in the graph denotes an atom and each edge denotes a bond. The two classes in NCI109 are about compounds’ ability to inhibit or interact with the specified cancer cell line [53].
ENZYMES is developed to classify each enzyme into one of six EC (Enzyme Commission) top-level classes. Each graph at hand describes a protein in a way that nodes represent secondary structure elements (SSEs) while edges show the relationship between these SSEs either in space or sequence [56].
@@
BZR consists of molecular graphs that represent chemical compounds. Each graph in this dataset is a chemical compound where nodes are atoms and edges are the chemical bonds between them. The classification task is focused on classifying compounds based on their biological activity with respect to binding with benzodiazepine receptors. In other words, the binary label of each graph can indicate binding or non-binding status to the benzodiazepine receptor [52].
@@
IMDB-M is a collection of social network graphs derived from the Internet Movie Database (IMDB). Each graph is about a collaboration network from movies whereby nodes stand for actors or actresses and edges indicate that the two actors appeared in the same movie. The dataset comprises three classes reflecting the movie genres [57].
COLORS-3 is a collection of artificially generated graphs. The dataset’s primary target is to classify these graphs based on their color distribution patterns and structural properties [58].
@@
D&D is a dataset of protein graphs. Each graph in this dataset represents a protein, with nodes corresponding to amino acids and edges corresponding to the spatial or sequential proximity between these amino acids. The primary objective of the D&D dataset is to classify proteins into one of two categories, i.e., enzymes or non-enzymes [55].
Table 7: Summary statistics of datasets for graph classification.
@@ D.2 Graph Regression
Table 8 provides an overview of the selected datasets in terms of their task numbers, compounds and other features, including splits, and metrics. A more detailed description is provided below.
QM8 is a benchmark dataset in computational chemistry, which is developed to facilitate the development and evaluation of machine learning approaches for quantum mechanical property prediction. It contains approximately 21,786 molecular structures, each characterized by their calculated properties using quantum chemistry methods, specifically focusing on electronic spectra [10, 87].
BACE is a collection of biochemical data used to evaluate computational methods for drug discovery. The dataset includes a total of 1,522 compounds, each annotated with their binding affinities, as well as molecular descriptors and fingerprints to facilitate the development and assessment of machine learning models [10, 88].
ESOL is a prominent resource in cheminformatics, which is leveraged to evaluate machine learning models on the prediction of aqueous solubility of small molecules. The dataset, derived from the work of Delaney, consists of a diverse range of chemical compounds with experimentally determined solubility values expressed in logS, where S is the solubility in mols per liter. It includes 1128 compounds, serving as a benchmark dataset for solubility prediction tasks [89, 10].
FreeSolv is a dataset containing hydration-free energies for small molecules in an aqueous solution. It comprises data for a wide range of organic molecules, providing both experimental values and calculated predictions based on molecular dynamics simulations [90, 10].
@@
Table 8: Details of datasets for graph regression.
@@ D.3 Node Classification
Table 9 presents descriptive statistics of the seven datasets used for node classification. It is evident that there is a significant variance in the scale of the selected datasets and each of them show different characteristics. Further background information and details about these datasets are provided below.
Cora comprises a collection of 2,708 scientific publications classified into seven distinct categories. Each publication in the dataset is represented as a node in a citation network, where edges indicate citation relationships between papers [59].
CiteSeer is a widely used citation network dataset. It comprises scientific publications categorized into six classes, with each publication represented by a 3,327-dimensional binary vector recording the presence or absence of specific words [59].
PubMed consists of scientific publications from the PubMed database, categorized into three classes based on their Medical Subject Headings (MeSH) terms. Each node is represented by a sparse bag-of-words vector derived from the content of the corresponding publication [59].
Table 9: Summary statistics of datasets for node classification.
@@
Cornell, Texas, and Wisconsin consist of nodes that represent web pages and edges which denote hyperlinks between these pages. Each node has a class label which typically denotes the topic of the web page; this allows tasks including node classification and link prediction to be performed. The datasets have different sizes. In particular, Cornell and Texas have 183 nodes while Wisconsin has 251 nodes. These networks are relatively small and sparse when compared to other larger citation datasets [60].
Github includes rich node attributes representing the diverse features of developers, such as their interests, skills, and contributions to various repositories. Additionally, the edges within the network capture the interactions and collaborations among developers, which result in a multi-faceted graph structure [61].
@@ D.4 Out-of-distribution shifts
Size shifts. For the selected datasets including D&D, PROTEINS, NCI1, and NCI109, we utilize the data provided by the previous work [66]. In this setup, the graphs with the smallest 50% of nodes are used as the training set, those with the largest 20% of nodes are used as the test set, and the remaining graphs are used as the validation set. The data can be downloaded from https://www.dropbox.com/s/38eg3twe4dd1hbt/data.zip.
Density shifts. For the selected datasets, we divide the datasets based on graph density. In particular, the 50% of graphs with the lowest density are used as the training set, the 20% with the highest density are used as the test set, and the remaining graphs are used as the validation set. After applying density shifts, the following densities are observed: for D&D, the training set density is 0.0274, the validation set density is 0.0567, and the test set density is 0.1245; for PROTEINS, the training set density is 0.1709, the validation set density is 0.4183, and the test set density is 1.0752; for NCI1, the training set density is 0.1229, the validation set density is 0.1976, and the test set density is 0.2922; for NCI109, the training set density is 0.1248, the validation set density is 0.2000, and the test set density is 0.2916.
@@ Appendix E Additional Experiments
Table 10 presents the results of fixing the backbone as GraphConv and setting the pooling ratios of the sparse pooling method to 0.1, 0.3, 0.5, 0.7, and 0.9, respectively. The pooling ratios for dense pooling remain unchanged. As shown in Table 10, with the increase in pooling ratio, the accuracy exhibits a fluctuating upward trend. Among the sparse pooling methods, HGPSLPool and KMISPool consistently demonstrate superior performance. However, changes in the pooling ratio also result in varying stability across different methods. Furthermore, compared to binary classification tasks, the accuracy improvement with the increasing pooling ratio is more evident in multi-class classification tasks. This indicate that researchers should consider using larger pooling ratios when conducting studies on multi-class classification.
@@
Table 11 presents the results of using GCNConv as the backbone model while varying the pooling ratio. It can be observed that the overall accuracy of GCNConv is lower than that of GraphConv. Generally, as the pooling ratio increases, classification accuracy improves and dense pooling methods are more powerful than sparse pooling methods at the cost of increased computation time. With the backbone of GCNConv, CGIPool and KMISPool perform best among the sparse pooling methods. Additionally, the impact of the pooling ratio on multi-class classification accuracy remains significantly greater than its impact on binary classification accuracy.
@@
Table 12 presents the results of varying the pooling ratio (0.1, 0.3, 0.5, 0.7, 0.9) in node classification tasks using GraphConv as the backbone model. Overall, the change in pooling ratio has a minimal impact on the accuracy of node classification tasks. For most methods, accuracy does not improve with an increased pooling ratio but rather fluctuates. The two best-performing methods overall are TopKPool and GSAPool. Additionally, higher pooling ratios result in increased computational costs for some methods, such as HGPSLPool and ASAPool. This indicates that when performing node classification, graph pooling methods should be used more cautiously, and excessive concern over pooling ratios may be unnecessary.
@@
Table 13 presents the results of node classification with GCNConv as the backbone model, varying the pooling ratio from 0.1, 0.3, 0.5, 0.7, to 0.9. It can be observed that, except for HGPSLPool, the performance of most methods with GCNConv is inferior to that with GraphConv. The performance of different pooling methods varies across different pooling ratios, making it difficult to identify a single pooling method that consistently outperforms others at all pooling ratios. Nevertheless, CGIPool and KMISPool still relatively demonstrate the best overall performance. Additionally, increasing the pooling ratio does not consistently improve node classification tasks. Lastly, GCNConv has higher memory consumption compared to GraphConv, particularly for ASAPool, PANPool, COPool, and HGPSLPool.
@@
In summary, based on the aforementioned experiments, GraphConv is recommended as the backbone model over GCNConv for both graph classification and node classification tasks due to its higher accuracy and lower computational resource consumption. In terms of graph classification tasks, a larger pooling ratio generally improves results, especially for multi-class classification tasks. In contrast, the choice of pooling ratio is not very important for node classification tasks.
@@
Table 10: Results of graph classification using GraphConv as backbone for different graph pooling methods and pooling ratios. $\pm$ denotes standard deviation. OOM denotes out of GPU memory. OOT denotes cannot be computed within 12 hours.
@@
Table 11: Results of graph classification using GCNConv as backbone for different graph pooling methods and pooling ratios. $\pm$ denotes standard deviation. OOM denotes out of GPU memory. OOT denotes cannot be computed within 12 hours.
@@
Table 12: Results of node classification using GraphConv as backbone for different graph pooling methods and pooling ratios. $\pm$ denotes standard deviation. OOM denotes out of GPU memory. OOT denotes cannot be computed within 12 hours.
@@
Table 13: Results of node classification using GCNConv as the backbone for different graph pooling methods and pooling ratios. $\pm$ denotes standard deviation. OOM denotes out of GPU memory. OOT denotes cannot be computed within 12 hours.